# Finetuning Large Language Model as an Effective Symbolic Regressor

### v1

@@ Introduction
For a SR dataset $D=\{(d_{i},y_{i})\}_{i=1}^{N_{\text{train}}}$, we first formulate the data matrix $d_{i}$ as a prompt $P_{i}$ involves two components: an data-shared instruction part $I$ and a data-specific value part $V_{i}$. The instruction $I$ consists of: 1). The definition of the SR task that requests the LLM to derive a symbolic equation from the provided data and a corresponding example, e.g., data to ’y = 2.0*sin(x) + 3.0’. 2). Basic operation scheme introduction, like existing commonly used symbolic operations, e.g., the definition of ’+’ and ’-’. The value part $V_{i}$ is formed by reformatting the $d_{i}$ as input-output pairs $\{(x_{i},y_{i})\}_{i=1}^{N}$ into a key-value structure (e.g., x_0=…, x_1=…, y=…) where $x_{i}\in\mathcal{R}^{K\times C_{i}}$ is the independent variable values and $y\in\mathcal{R}^{K}$ is the dependent one.
@@
The prompt $P_{i}$ is fed into the LLM backbone, which in turn generates the predicted equation $\hat{f}_{i}$}. Similar to most language models, we employ cross-entropy loss which constrains the similarity between estimated and ground-truth tokens, which can be presented as:
@@
For parameter-efficient fine-tuning, we employ Low-Rank Adaptation (LoRA)  (Hu et al. 2022) on the LLM backbone.
Reinforcement Tuning by Form-GRPO
@@
Numerical Reward We use the numerical reward $\mathcal{R}_{\text{numerical}}$ to quantify how well an equation fits the input-output data. This reward is only utilized for equations passed the check of the valid decision function. We use the $R^{2}$ score to achieve the numerical reward. However, since the raw $R^{2}$ score is unbounded from below, its direct use can lead to training instability. We therefore apply a truncation strategy, defining the final reward as follows:
@@
Equivalent Reward We define the equivalence reward $\mathcal{R}_{\text{equiv}}$ to incentivize the generated equations that have a completely equivalent form to the ground truths (do not need the same coefficients). Given the difficulty and critical importance of achieving a perfect structural match, we apply a significant bonus for this accomplishment. The reward is computed as follows:
@@
To leverage this composite reward signal for policy optimization, we employ the Group Relative Policy Optimization (GRPO) algorithm (Shao et al. 2024). Specifically, for a given input prompt, the process begins by sampling a group of $G$ candidate equations from the reference policy, where $G$ is set as 8 following the GRPO original setting. $\pi_{\theta_{\text{old}}}$. Each equation is then evaluated to obtain a corresponding set of rewards, which are subsequently normalized using the group’s mean and standard deviation. This single normalized reward serves as the advantage estimate for all tokens within the corresponding equation.
Scientific Generative Agent Refinement
Following the initial stages of Instruction Tuning and Form-GRPO, our LLM is equipped with robust capabilities for symbolic regression. To unlock its full potential, we employ Scientific Generative Agent Refinement, an iterative strategy designed to progressively generate and refine equations (Ma et al. 2024). Scientific Generative Agent leverages the LLM as a core generative engine within a structured search process. The procedure unfolds as follows:
Initialization The process is initialized by prompting the LLM to generate a single primitive equation. This equation is then evaluated to compute its initial score, and this first individual seeds the population.
Iterative Refinement Cycle In subsequent iterations, the topk-scoring equations are identified and embedded into a new prompt, using them as in-context examples to steer the model’s generation process. The subsequent creation of ”offspring” is carefully controlled by varying the sampling temperature, which balances the refinement of existing solutions (exploitation) with the discovery of novel ones (exploration). This allows the agent to systematically navigate the solution space towards an optimal equation.
Evaluation and Population Update Each newly generated offspring is evaluated to determine its fitness, quantified by the $R^{2}$. To enhance precision, a conditional optimization step refines the equation’s constants before the evaluated offspring and its score are integrated back into the population for the next cycle.
Termination After completing the specified number of iterations, the cycle terminates. The final solution is determined by selecting the single equation with the best score from the entire population.
Experiments
Methods Type $S_{\text{struct}}$ (gpt-4o) $S_{\text{struct}}$ (rule) $R^{2}$ $\text{Acc}_{\tau}$ Time Cost (s)
gplearn GP 0.266 0.296 0.2 0.074 14.05
AFP GP 0.264 0.371 0.245 0.06 3.74
AFP-FE GP 0.298 0.389 0.296 0.072 4.12
GP-GOMEA GP 0.337 0.356 0.36 0.238 11.32
uDSR GP 0.244 0.422 0.349 0.027 237.89
PySR GP 0.368 0.382 0.663 0.398 6.85
LLM-SR(gpt-3.5-turbo) LLM-based 0.291 0.437 0.255 0.189 334.75
LLM-SR(gpt-4o-mini) LLM-based 0.299 0.415 0.355 0.221 413.27
LLM-SR(Qwen2.5-7B) LLM-based 0.306 0.404 0.313 0.205 465.94
SGA(gpt-3.5-turbo) LLM-based 0.342 0.504 0.327 0.17 57.47
SGA(gpt-4o-mini) LLM-based 0.351 0.492 0.286 0.168 69.12
SGA(Qwen2.5-7B) LLM-based 0.362 0.480 0.273 0.176 73.81
SymbolicChat LLM-based 0.429 0.602 0.799 0.412 92.05
@@
LLM Backbone Ins. tuning Re. tuning Iter Stratgy $S_{\text{struct}}$ (gpt-4o) $S_{\text{struct}}$ (rule) $R^{2}$ $\text{Acc}_{\tau}$
Qwen2.5-7B 0.353 0.466 0.201 0.145
@@
Qwen2.5-7B-SFT-GPRO 0.403 0.574 0.540 0.252
LLM-SR 0.426 0.592 0.705 0.394
SGA 0.429 0.602 0.799 0.412
Table 3: Ablation Study of SymbolicChat. The best and second-best figures are in bold and underlined, respectively.
Implementation Details
@@
We employ Qwen2.5-7B-Instruct (Team 2024) as our LLM backbone. For constructing all prompts and evaluating baseline methods, we consistently sample 200 input-output $(x,y)$ pairs from the dataset. During the instruction Tuning stage, we utilize a dataset of 146,590 unique expressions, creating five distinct prompt variations for each, and fine-tune the model using LoRA with a rank of 8 and an alpha of 32. For the rule-based reward Tuning stage, we also use a set of 1,000 expressions and sample 8 candidate responses from the LLM for each input. We configure the reward function hyperparameters to 1.0, 2.0, 2.0, and 4.0, with justification provided in the Appendix. All other settings are kept identical to the preceding stage. Finally, in the scientific generative agent refinement stage, we configure the agent to run for 5 iterations, generating 2 expressions for exploitation and 4 for exploration within each cycle. More implementation and training details are provided in the supplementary material.
Comparison with state-of-the-art methods.
@@
As detailed in Table 2, our proposed method, SymbolicChat, outperforms existing baselines across most key evaluation metrics. Existing LLM-based SR methods often failed to compare with traditional methods, proving the barrier of treating complex equations introduced by the gap between the pre-trained smooth-oriented knowledge and the SR’s high accuracy demands. Fortunately, our SymbolicChat suppresses the traditional state-of-the-art PySR (Cranmer 2023) on structure metrics and $\mathcal{R}^{2}$ by a large margin, also achieving comparable results on $\text{Acc}_{\tau}$, proving the effectiveness of SymbolicChat and the representativeness of the sample in our SymbArena that utilizes only 10 thousands data lead to new state of the art. Compared with the previous state-of-the-art LLM-based SR method, LLM-SR (Shojaee et al. 2024), our model achieves at least 0.1 structure increases and brings about 2-fold $R^{2}$ gains, only with one-fourth of the inference time cost. Also, compared with a state-of-the-art LLM science discovery model, SGA, our methods provide more advantages, further demonstrating the crucial importance of the fine-tuning for enhancing the symbolic regression ability of LLM in SR tasks.
Ablation study
@@
To systematically evaluate the contributions of each core component in the SymbolicChat framework, we conducted a series of detailed ablation studies. These studies were designed to quantify the independent effectiveness of instruction tuning, rule-based reward tuning, and the scientific generative agent’s iterative strategy. Our analysis begins with the untuned Qwen2.5-7B backbone, which serves as a baseline by achieving an $R^{2}$ score of 0.201. This initial result indicates that a general-purpose LLM, despite its foundational capabilities, is insufficient for high-precision symbolic regression. Furthermore, directly applying iterative strategies like SGA to this base model did not yield effective performance improvements.
@@
With the introduction of instruction tuning, the model showed significant improvement, with its $R^{2}$ score increasing to 0.313 and its rule-based form-level consistency score rising substantially from 0.466 to 0.590. This demonstrates the critical role of this step in adapting the model to the task’s format and enabling it to generate structurally correct expressions. Building on this foundation, we introduced rule-based reward tuning using the GRPO algorithm, which led to a substantial leap in performance, boosting the $R^{2}$ score from 0.201 to 0.540. This provides strong evidence that our designed reward mechanism effectively guides the model to generate expressions that are not only structurally sound but also numerically fit the given data with high fidelity. Finally, after applying the scientific generative agent iterative strategy, the complete SymbolicChat model reached its peak performance, with the $R^{2}$ score further increasing from 0.540 to 0.799. This shows that using the fine-tuned LLM as a core generator and meticulously refining solutions through an iterative search of the solution space is the decisive step in achieving optimal numerical fitting and elevating the model’s performance to a new state-of-the-art.
Conclusion
@@
In this work, we delve into the limitations of the effectiveness and accuracy of LLM-based Symbolic Regression (SR) methods as the lack of SR-specific fine-tuning. To tackle this, we propose a SymbArena specifically designed for LLM fine-tuning, associated with a novel evaluation metric beyond traditional numerical and covering form-level consistency. With this, we propose a new LLM-based SR strong baseline, SymbolicChat, achieving a new state of the art of LLM-based SR and suppressing traditional SR methods. Through these efforts, we hope to make modest contributions to the field of LLM scientific discovery, fostering the development of more robust and adaptable tools for future symbolic regression.
Supplementary Material
Implementation Details
Experimental Environment Setup
@@
Table 4: Experimental Environment Setup.
Traditional SR Methods
We compare our SymbolicChat against several state-of-the-art traditional Symbolic Regression (SR) baselines, including gplearn, AFP (Schmidt and Lipson 2010), AFP-FE (Schmidt and Lipson 2009), GP-GOMEA (Virgolin et al. 2021), uDSR (Landajuela et al. 2022), and PySR (Cranmer 2023). The hyperparameters for all baseline methods are set to the default values as specified in their original publications. A detailed summary of these parameters is provided in Table 5:
@@
Table 5: Hyper-parameter setting of traditional SR methods.
LLM-Based SR Methods
@@
2) Search Prompt (using in SymbolicChat, LLM-SR and SGA): {graybox} You can also refer to the previously proposed formula and their fitness(the smaller the better) stored in {pred_dict}:
  $0:[3*sin(x_{0})+x_{1}+2.4,\hskip 5.0pt0.01]$

### v2

@@ 3.1 Instruction Tuning
For a SR dataset $D=\{(d_{i},f_{i})\}_{i=1}^{N_{\text{train}}}$, we first formulate the data matrix $d_{i}$ as a input representation $P_{i}$ involves two components: an data-shared instruction part $I$ and a data-specific value part $V_{i}$. The instruction $I$ consists of: 1). The definition of the SR task that requests the LLM to derive a symbolic equation from the provided data and a corresponding example, e.g., data to ’f(x) = 2.0*sin(x) + 3.0’. 2). Basic operation scheme introduction, like existing commonly used symbolic operations, e.g., the definition of ’+’ and ’-’. The value part $V_{i}$ is formed by reformatting the $d_{i}$ as input-output pairs $\{(x_{i},f(x_{i}))\}_{i=1}^{K}$ into a key-value structure (e.g., x_0=…, x_1=…, f(x)=…) where $x_{i}\in\mathcal{R}^{K\times C_{i}}$ is the independent variable values and $f(x_{i})\in\mathcal{R}^{K}$ is the dependent one.
@@
The input $P_{i}$ is fed into the LLM backbone, which in turn generates the predicted equation $\hat{f}_{i}$}. Similar to most language models, we employ cross-entropy loss which constrains the similarity between estimated and ground-truth tokens, which can be presented as:
@@
For parameter-efficient fine-tuning, we employ Low-Rank Adaptation (LoRA)  (Hu et al., 2022) on the LLM backbone.
3.2 Reinforcement Tuning by Form-GRPO
@@
Numerical Reward. We use the numerical reward $\mathcal{R}_{\text{numerical}}$ to quantify how well an equation fits the input-output data. This reward is only utilized for equations passed the check of the valid decision function. We use the $R^{2}$ score to achieve the numerical reward. However, since the raw $R^{2}$ score is unbounded from below, its direct use can lead to training instability. We therefore apply a truncation strategy, defining the final reward as follows:
@@
Equivalent Reward. We define the equivalence reward $\mathcal{R}_{\text{equiv}}$ to incentivize the generated equations that have a completely equivalent form to the ground truths (do not need the same coefficients). Given the difficulty and critical importance of achieving a perfect structural match, we apply a significant bonus for this accomplishment. The reward is computed as follows:
@@
To leverage this composite reward signal for policy optimization, we employ the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024). Specifically, for a given input, the process begins by sampling a group of $G$ candidate equations from the reference policy, where $G$ is set as 8 following the GRPO original setting. Each equation is then evaluated to obtain rewards, which are subsequently normalized using the group’s mean and standard deviation. This single normalized reward serves as the advantage estimate for all tokens within the corresponding equation.
3.3 Hypothesis–Experiment–Revision inference
During inference, we introduce a Hypothesis–Experiment–Revision (HER) framework, an iterative strategy designed to yield more reliable equations. The core idea is to generate multiple candidate hypotheses in each iteration, validate them through quantitative experiments, and incorporate reflective revision to selectively retain high-quality candidates—thereby emulating the scientific cycle of hypothesis formulation, experimental validation, and reflective refinement.
Hypothesis: Given an input $P$, the Symbolic-R1 model produces a set of candidate equations $\{\hat{f}_{1},\hat{f}_{2},\ldots,\hat{f}_{K}\}$, where $K$ is the number of hypotheses (set to 6, following prior multi-hypothesis discovery approaches such as SGA (Ma et al., 2024)). For each candidate, we employ numerical optimization tools to refine the coefficients and improve predictive accuracy.
Experiment: Each hypothesis is then quantitatively evaluated using the $R^{2}$ score, which serves as an analogue to experimental validation of the consistency between a hypothesis and real-world observations.
Revision: The experimental outcomes are subsequently fed back into Symbolic-R1, prompting it to generate qualitative assessments that resemble scientists’ reflective notes on each hypothesis. These evaluations are jointly stored as tuples $(equation,score,comments)$ in a memory bank. The bank maintains the top-5 records ranked by $R^{2}$, updating in each iteration by discarding weaker candidates and preserving the most promising ones.
The curated memory is then reformulated into a reflection prompt and appended to the input $P$, enabling Symbolic-R1 to generate refined hypotheses in the next iteration. This iterative loop gradually improves the reliability and scientific plausibility of the inferred equations.
4 Experiments
@@
Methods Type $S_{\text{struct}}$ (gpt-4o) $S_{\text{struct}}$ (rule) $R^{2}$ $\text{Acc}_{\tau}$
gplearn GP $0.266$ $0.296$ $0.200$ $0.074$
AFP GP $0.264$ $0.371$ $0.245$ $0.060$
AFP-FE GP $0.298$ $0.389$ $0.296$ $0.072$
GP-GOMEA GP $0.337$ $0.356$ $0.360$ $0.238$
uDSR GP $0.244$ $0.422$ $0.349$ $0.027$
PySR GP 0.368 $0.382$ 0.663 0.398
LLM-SR(gpt-3.5-turbo) LLM-based $0.291$ $0.437$ $0.255$ $0.189$
LLM-SR(gpt-4o-mini) LLM-based $0.299$ $0.415$ $0.355$ $0.221$
LLM-SR(Qwen2.5-7B) LLM-based $0.306$ $0.404$ $0.313$ $0.205$
SGA(gpt-3.5-turbo) LLM-based $0.342$ 0.504 $0.327$ $0.170$
SGA(gpt-4o-mini) LLM-based $0.351$ $0.492$ $0.286$ $0.168$
SGA(Qwen2.5-7B) LLM-based $0.362$ $0.480$ $0.273$ $0.176$
Symbolic-R1 LLM-based $0.436$ $0.607$ $0.808$ $0.404$
Table 3: Ablation Study of Symbolic-R1. This table evaluates the individual contribution of each core component: instruction tuning (IFT), reinforcement tuning (RFT), and the Hypothesis–Experiment–Revision (HER) framework. The best and second-best figures are in bold and underlined, respectively.
LLM Backbone Iteration Strategy $S_{\text{struct}}$ (gpt-4o) $S_{\text{struct}}$ (rule) $R^{2}$ $\text{Acc}_{\tau}$
Qwen2.5-7B None 0.353 0.466 0.201 0.145
@@
HER 0.456 0.613 0.624 0.309
  +IFT +RFT None 0.403 0.574 0.540 0.252
LLM-SR 0.396 0.538 0.743 0.343
SGA 0.410 0.569 0.778 0.373
HER 0.436 0.607 0.808 0.404
4.1 Implementation Details
@@
We employ Qwen2.5-7B-Instruct (Team, 2024) as our LLM backbone. For constructing all prompts and evaluating baseline methods, we consistently sample 200 input-output $(x,y)$ pairs from the dataset. During the instruction tuning stage, we utilize a dataset of 146,590 unique expressions, creating five distinct prompt variations for each, and fine-tune the model using LoRA with a rank of 8 and an alpha of 32. For the reinforcement tuning stage, we also use a set of 1,000 expressions and sample 8 candidate responses from the LLM for each input. We configure the reward function hyperparameters to 1.0, 2.0, 2.0, and 4.0, with justification provided in the Appendix. All other settings are kept identical to the preceding stage. Finally, in the Hypothesis–Experiment–Revision inference stage, we configure the llm to run for 5 iterations, generating 6 equations within each cycle. More implementation and training details are provided in the supplementary material.
4.2 Comparison with state-of-the-art methods.
@@
As detailed in Table 2, our proposed method, Symbolic-R1, outperforms existing baselines across most key evaluation metrics. Existing LLM-based SR methods often failed to compare with traditional methods, proving the barrier of treating complex equations introduced by the gap between the pre-trained smooth-oriented knowledge and the SR’s high accuracy demands. Fortunately, our Symbolic-R1 suppresses the traditional state-of-the-art PySR Cranmer (2023) on structure metrics and $\mathcal{R}^{2}$ by a large margin, also achieving comparable results on $\text{Acc}_{\tau}$, proving the effectiveness of Symbolic-R1 Compared with the previous state-of-the-art LLM-based SR method, LLM-SR Shojaee et al. (2024), our model achieves at least 0.1 structure increases and brings about 2-fold $R^{2}$ gains, only with one-fourth of the inference time cost. Also, compared with a state-of-the-art LLM science discovery model, SGA, our methods provide more advantages, further demonstrating the crucial importance of the fine-tuning for enhancing the symbolic regression ability of LLM in SR tasks.
4.3 Ablation study
@@
To systematically evaluate the contributions of each core component in the Symbolic-R1 framework, we conducted a series of detailed ablation studies. These studies were designed to quantify the independent effectiveness of instruction tuning, reinforcement tuning, and the Hypothesis–Experiment–Revision framework. Our analysis begins with the untuned Qwen2.5-7B backbone, which serves as a baseline by achieving an $R^{2}$ score of 0.201. This initial result indicates that a general-purpose LLM, despite its foundational capabilities, is insufficient for high-precision symbolic regression. Furthermore, directly applying iterative strategies like SGA to this base model did not yield effective performance improvements.
@@
With the introduction of instruction tuning, the model showed significant improvement, with its $R^{2}$ score increasing to 0.313 and its rule-based form-level consistency score rising substantially from 0.466 to 0.590. This demonstrates the critical role of this step in adapting the model to the task’s format and enabling it to generate structurally correct expressions. Building on this foundation, we introduced reinforcement tuning using the GRPO algorithm, which led to a substantial leap in performance, boosting the $R^{2}$ score from 0.201 to 0.540. This provides strong evidence that our designed reward mechanism effectively guides the model to generate expressions that are not only structurally sound but also numerically fit the given data with high fidelity. Finally, after applying the Hypothesis–Experiment–Revision framework, the complete Symbolic-R1 model reached its peak performance, with the $R^{2}$ score further increasing from 0.540 to 0.808. This shows that using the fine-tuned LLM as a core generator and meticulously refining solutions through an iterative search of the solution space is the decisive step in achieving optimal numerical fitting and elevating the model’s performance to a new state-of-the-art.
5 Conclusion
@@
In this work, we delve into the limitations of the effectiveness and accuracy of LLM-based symbolic regression methods (see Appendix A for a full review) as the lack of SR-specific fine-tuning. To tackle this, we propose a SymbArena specifically designed for LLM fine-tuning, associated with a novel evaluation metric beyond traditional numerical and covering form-level consistency. With this, we propose a new LLM-based SR strong baseline, Symbolic-R1, achieving a new state of the art of LLM-based SR and suppressing traditional SR methods. Through these efforts, we hope to make modest contributions to the field of LLM scientific discovery, fostering the development of more robust and adaptable tools for future symbolic regression.
Appendix
Appendix A Related work
A.1 Traditional Symbolic Regression
Symbolic regression is a machine learning technique for discovering concise, interpretable mathematical equations from data (Wang et al., 2019). Early approaches were dominated by Genetic Programming (GP), which proved effective in identifying physical laws from experimental data (Schmidt & Lipson, 2009). To overcome the efficiency and scalability limitations of GP, methods based on sparse regression were developed. A prominent example is the Sparse Identification of Nonlinear Dynamics (SINDy) algorithm, which identifies governing equations by applying sparse optimization to a library of candidate functions (Brunton et al., 2016; Rudy et al., 2017). More recently, deep learning has catalyzed major advances across the entire SR workflow. For instance, researchers use reinforcement learning to frame equation generation as a sequential decision problem (Li et al., 2025b; Xu et al., 2024) or apply pre-trained transformers to rapidly produce candidate equation structures (Biggio et al., 2021; Kamienny et al., 2022; Shojaee et al., 2023; Valipour et al., 2021; Vastl et al., 2024). Deep learning also enhances foundational aspects; autoencoders can learn low-dimensional latent data representations, facilitating the discovery of simpler models in an optimal coordinate system (Champion et al., 2019; Li et al., 2025a). Furthermore, hybrid methods show great promise, such as using RNNs to generate high-quality initial populations for more efficient genetic programming searches (Mundhenk et al., 2021; Landajuela et al., 2022).
A.2 LLM for Symbolic Regression
Recent SR methods leverage LLMs as hypothesis generators, iteratively refining equations with data-driven feedback to embed physical knowledge (Li et al., 2025a). LLM-SR (Shojaee et al., 2024) uses the LLM as a black-box optimizer for self-improvement or as an evolutionary engine performing crossover and mutation. SGA (Ma et al., 2024) employs a bilevel optimization where the LLM generates discrete equation structures (upper level) and a differentiable simulator optimizes their continuous parameters (lower level). In contrast, LaSR (Grayeli et al., 2024) uses an LLM to evolve a library of abstract textual concepts that guide and accelerate a separate evolutionary search. However, a recent benchmark (Shojaee et al., 2025) reveals a key limitation: the performance of general-purpose LLMs plummets when they are prevented from reciting known equations. This highlights the need for domain-specific model adaptation over relying on pure inference.
Appendix B Implementation Details
B.1 Experimental Environment Setup
@@
B.2 Traditional SR Methods
We compare our Symbolic-R1 against several state-of-the-art traditional Symbolic Regression (SR) baselines, including gplearn, AFP (Schmidt & Lipson, 2010), AFP-FE (Schmidt & Lipson, 2009), GP-GOMEA (Virgolin et al., 2021), uDSR (Landajuela et al., 2022), and PySR (Cranmer, 2023). The hyperparameters for all baseline methods are set to the default values as specified in their original publications. A detailed summary of these parameters is provided in Tab. 5:
Table 5: Hyper-parameter setting of traditional SR methods.
@@ B.2 Traditional SR Methods
B.3 LLM-Based SR Methods
@@ B.4.2 Message
2) Memory Prompt (using in Symbolic-R1): {graybox} You can refer to the previously proposed formulas and their corresponding fitness scores (lower is better), which are stored in pred_dict:
  • $0:$
    $3\sin(x_{0})+x_{1}+2.4,\quad 0.01]$
Based on the analysis, here are some suggestions for improvement:
  • Consider adding a term that captures the interaction between $x_{0}$ and $x_{1}$, such as $0.479-0.476x_{1}-0.231x_{0}$, to refine the model.
  • Introduce a quadratic term for $x_{0}$ to capture non-linear effects, for example, $0.21-0.77x_{0}^{2}$.
Please consider these suggestions when generating new formulas.